OpenStudy (anonymous):

What is the distance between points (12, 4) and (18, –1)? Round to the nearest tenth.

OpenStudy (anonymous):

\[\sqrt{(18-12)^2+(4+1)^2}=\sqrt{36+25}=\sqrt{61} \approx7.81\]